Given the following market function, find the equilibrium price and quantity:

 [6 marks]

Qd = 54 – 4p

Qs = 4 + 4p



1
Expert's answer
2021-12-20T10:55:26-0500

At equilibrium,

"Q_d=Q_s"

"54-4p=4+4p"

"54-4=4p+4p"

"50=8p"

"P=6.25"

"Q_d=54-4p=54-4(6.25)=29."

"\therefore" Equilibrium price and quantity are 6.25 and 29 respectively.